327 MHz Gregorian Dome Receiver Log

as of Monday, 18-Dec-2017 12:35:15 AST

Go to Most Recent Entry


NameEmailDateComment
ANS anolla@naic.edu 8/4/2000 These receiver was installed during July , 2000.It was calibrated on 8/3/2000 ,the cal values are Chn. A 186 k deg & Chn B is a 150 K deg.
ANS anolla@naic.edu 8/25/2000 Dewar taken down for repairs, it is not holding the vacuum.
ANS anolla@naic.edu 9/20/2000 Dewar was taken up on 9/18/2000.It was tested in the cryolab for more than a week. Coco found a couple of leaks at the Dewar; also he replaced the crosshead on it. New Cal & RF cables were done
ANS anolla@naic.edu 10/10/2000 Receiver warmed up due to power dip on Saturday that tripped off the compressor. Was cooled down.
ANS anolla@naic.edu 10/11/2000 There is a problem with the receiver cryo and it looks like if the compressor is not holding the 3 crossheads attached to it.
ANS anolla@naic.edu 11/8/2000 Start cooling down using the Austin small cryo comp.
ANS anolla@naic.edu 11/9/2000 Dewar cooled , cryo temp is 49.2 ºk. Austin comp 125 used.
ANS anolla@naic.edu 11/14/2000 Dewar at room temp due that was necessary to use the Austin 125 (small) cryo comp. to cool down the 430 Mhz Dewar.
ANS anolla@naic.edu 12/13/2000 Channel B RF out was not connected .ANS hooked up a 327 Mhz filter on it.
ANS anolla@naic.edu 12/15/2000 JR removed the matching xformers & hybrids out of the antenna.
ANS anolla@naic.edu 12/28/2000 JR move the receiver to the new open frame cabinet at the rotary floor.
ANS anolla@naic.edu 1/31/2001 Receiver cooled with comp. 008.
ANS anolla@naic.edu 2/8/2001 The A/L power supply was not working is suspected that the DS chip is burned out.
ANS anolla@naic.edu 3/12/2001 Dewar hot due to cryo comp shutdown, comp off since 3/10/2001.
ANS anolla@naic.edu 4/4/2001 Dewar turned off to replace crosshead. Also Dewar will be taken down to check for leaks.
ANS anolla@naic.edu 4/20/2001 Dewar was tested for leaks with the Varian leak tester, no leak was found. The crosshead needs to be cleaned out.
ANS anolla@naic.edu 5/23/2001 A/L sw at the radar room bypassed , always on antenna except when 430 mhz xmmter up at CH 1.
ANS anolla@naic.edu 7/27/2001 The recvr was installed back, but there is not ant. System with 50 ohms at input.
ANS anolla@naic.edu 8/15/2001 The 327 kHz antenna was installed; it has new hybrids on it ( TRM ) .Also the match xformers were removed from it.
ANS anolla@naic.edu 8/21/2001 The channel A A/L sw is bad . Also this morning due to a power glitch the cryo compressor tripped off.
ANS anolla@naic.edu 8/22/2001 System is cold. Just VP to tell, problems with the Lakeshore reading. JR is working with the A/L driver.
ANS anolla@naic.edu 8/23/2001 The A/L driver was ok , the problem was the 24 vdc power supply which drives it had a 7 v ripple that was causing the A/L relays to chattering.
ANS anolla@naic.edu 8/27/2001 The cryo comp replaced , it was ED 007 (back to WSB) now is ED 016. Also this morning the comp was found tripped off , maybe due to power dips.
ANS anolla@naic.edu 8/30/2001 The system has been cooling down because it warmed up this morning when we try to switch cryo compressors that didn't worked out , comp ED 011 still has problems.
ANS anolla@naic.edu 9/20/2001 System been cooled with compressor ED 021 ( Austin 125 ).It had a leak on it, but it is not in its cryo lines , we check them and they are ok.
ANS anolla@naic.edu 9/20/2001 System cooled. Need Tsys check out. Also the problem with its cryo reading fixed by itself , what it was ? we don't know.
ANS anolla@naic.edu 9/27/2001 System again hot , problems with the cryo. System at room temp.
ANS anolla@naic.edu 10/18/2001 Jose Rosa modified the Lakeshore ckt in order to be able to take downstairs readings . Also WG found a problem with the cryo reading board ( loose Vcc cable on it). Both chassis were installed back at the Dome.
ANS anolla@naic.edu 10/21/2001 327 mhz protected from the SB xmitter thru the CB/SBW shutter control relay ( if shutter is closed it will send the 327 mhz to load, so have to be aware of that )
ANS anolla@naic.edu 1/30/2002 Chan A amp is sn 350-12 and its biases are: First stage:Vd 1.42 vdc,Id 3.7 ma Second stage:Vd 1.44 vdc,Id 4.9 ma. Chan B amp is sn 350-14 and its biases are: First stage:Vd 1.51 vdc, Id 4.3 ma Second stage: Vd 1.51 vdc, Id 5.0 ma.
ANS anolla@naic.edu 4/1/2002 Spare amp 350-13 came fixed from NRAO.
ANS anolla@naic.edu 7/31/2002 New receiver at room temp installed.
LW lwray@naic.edu 8/5/02 2 MHz comb that was seen in the band has been traced to the radar box. The box has been re-shielded and re-installed in the dome. Data was taken today with the box ON. The 2 MHz comb is still there, but VERY much reduced in power.
ansanolla@naic.eduTuesday, October 29, 2002 11:32:29New Angle Linear amps on it , model 2X327GST (room temp).Channel A is S/N 11467 & Channel B is S/N 11468 .Spare is S/N 11469.They are 20º better than the former ones.
LSWllwray@naic.eduMonday, November 4, 2002 15:01:33Tested LNA box using NFA, found 65K with 5 dB ENR, and ~55K with 15 dB ENR. Tsys program reads 110K at 9am. (both channels read ~same in these cases) At 2pm, Tsys program reads 150K. RFI environment is the cause of the increase.
ansaanolla@naic edu.Friday, December 6, 2002 13:02:39Connectors added to the antenna cables to protect them from been broken.
lswlwray@naic.eduThursday, January 22, 2004 16:08:35Jan 7: receiver and feed moved.
lswlwray@naic.eduThursday, January 22, 2004 16:09:44Jan 21: feed hooked up (thanks to PP) and receiver functioning in its new spot on the rotary floor (next to X Band)
lswllwray@naic.eduThursday, July 1, 2004 15:08:19Calixto and Carlos testing for RFI today.
ansanolla@naic.eduThursday, September 30, 2004 13:34:56When installing the hybrid for the polarization test of the system we found sort a few things , first : channel A antenna connector was unsolder , several sma connectors loose at the amp chassis. The Tsys with system polarized were ; chann A =119 deg , chann B=120 deg.
ansanolla@naic.eduMonday, October 4, 2004 13:06:31Tsys was +/- 20 K deg higher at chann B than chann A. Leading cables from the antenna to the recvr were unsolder. After fixing the cables both channels were at 120 K deg
ansanolla@naic.eduMonday, October 11, 2004 12:38:16System been polarized with Quad Hybrid.
ansanolla@naic.eduTuesday, October 12, 2004 13:02:48System been kept polarized with quad hybrid.
testtestWednesday, March 30, 2005 14:41:02test
W. Genterwgenter@naic.eduWednesday, September 7, 2005 09:42:28There has been a 16KHz comb in the bandpass of this receiver since May 2005. This did not affect pulsar work, but for all others this comb prevented observations in this band. Several efforts were made to find the source of this without success. In late August 2005, Rey Velez, PP and Myself found the source of the interference. The source is the Alfa rotation control motor. The Motor control cable from to the rotation control motor was not grounded to the chassis. At this time, the Alfa rotation drive motor must have power removed to observe at 327MHz.
ansanolla@naic.eduTuesday, November 29, 2005 10:53:04The 300 Mhz HP filters for RFI didn't work as expected.No change.
ansanolla@Wednesday, November 30, 2005 13:27:20The 300 Mhz HP filters were removed.Didn't work properly.
ansanolla@naic.eduThursday, April 13, 2006 13:02:21The rf cable of channel A between the front end and the filter at the post amp was replaced , was bad. Also the cal cable of channel B between post amps and the antenna was replaced.
ansanolla@naic.eduMonday, May 15, 2006 14:45:30The missing NRAO cooled amps were found in a box at the cryo shop.Back to 11F-K04-1 drawer in the stock room.
W. Genterwgenter@naic.eduMonday, July 24, 2006 10:55:35The new Dewar is in process of being finished this date. The heat shields are being plated. Dewar (no paint) directional couplers have been received and the cryo amplifiers located.
W. Genterwgenter@naic.eduWednesday, August 16, 2006 07:43:18The new dewar has arrived.
W. Genterwgenter@naic.eduThursday, September 14, 2006 09:42:24Two DCI cavity filters have been received for this receiver. The filters are centered at 327MHz with a BW of 50MHz.
ansanolla@naic.eduFriday, November 17, 2006 11:35:17Post amp chassis down for the filter bank installation.System down until further notice.
ansanolla@naic.eduWednesday, November 22, 2006 16:11:50327 Mhz back to service.New filter bank on it , but can't swap filters yet.Set to 327Mhz / 30 Mhz B.W.
ansanolla@naic.eduWednesday, November 29, 2006 14:23:18Filter bank control installed.Is bit 402 of the HP3488 I. The 50 Mhz B.W. filter is the default one.
ansanolla@naic.eduMonday, December 11, 2006 12:51:07The receiver configuration was reset in order to have the shortest rf cables between the DCI cavity filters & the post amp.So the DCI filters are at the bottom of the rf cabinet with the post amp , the recvr control was removed to the middle of the rf cab (open frame).
ansanolla@naic.eduFriday, December 22, 2006 08:59:14Due to the new configuration with the filter bank , the system gain was increased by 6 db when using the 50 Mhz B.W. filters .So a fix 6 db pad was added to the output of the rf of it to reduce these gain.Gain came from 3 db from doubling the B.W. and 3 more db from less loss of the 50 Mhz B.W. filter.
ansanolla@naic.eduMonday, April 23, 2007 09:53:04System down due to painting project.
crvFriday, June 6, 2008 10:45:25On friday May the 30th the new dewar was installed and the chasis were relocated.
crvFriday, June 6, 2008 10:47:49Tuesday June the 3rd the dewar amp were tested an the cryo temp checked finding problems with Cha Stage 2 amp ccurrent.
crvFriday, June 6, 2008 10:50:05Wednesday june the 4th the dewar was open and the problem was repaired. power connector over thighten. The connector was lossen and thight corectly.
crvTuesday, June 10, 2008 10:24:17Today the new dewar as started cooldown with a small compresssor ed 021.
crvWednesday, June 18, 2008 16:00:40On friday june the 12th the reciever being used is the coolreciever.
crvWednesday, June 18, 2008 16:04:42 50 MHz filter have been mounted over the feed horn and attached under the floor in brackets.
CRVMonday, October 13, 2008 14:16:11On friday october the 10th the crosshead stop rotating. Motor was stucked in one position. Crosshead ED 20 was removed Large amount of water found on both stage displacers. Dewar cylinder was cleaned and dried and a brand new frig was installed.
Stephen Jensensjensen@naic.eduThursday, December 11, 2008 16:32:494:00 PM: The 327 receiver is warming up nicely.
crvThursday, August 6, 2009 15:10:26reciever was vacuumed and is cooled down today 8/6/2009 after sunday night power failure and wednesday power failure .
CRVMonday, January 4, 2010 16:46:51System was removed from Comp. ED015 and connected to ALFA'S new compressor C2 . This compressor has now three cryo pumps (alfa 2, 327MHz , XB).
crvWednesday, January 13, 2010 10:09:32System was warming up on 1/12/10 70k = 165k. The system was totally warmed up vac.pump connected and will be coolled on 1/13/10.
crv Monday, February 22, 2010 14:16:38system is now connected to ALFA C1 compressor.
criosFriday, April 30, 2010 17:24:02system was disconnected from Compressor C2 and conected on the spare comp. ED015 to allow conecting Alfa second cryo pump to the C1 compressor. This change was made due to C2 comp. failing.
crios Friday, April 30, 2010 17:24:58The change of compressor was on April 23rd
crvWednesday, June 2, 2010 10:41:14The system is down due to an internal broken line on the compressor C3. Below the main dish is still flooded and is dificult to lower or raise a compressor.
crvMonday, September 27, 2010 17:02:59system was warmed and purged an recooled is operational.
GRgrivera@naic.eduTuesday, November 30, 2010 16:22:37The dewar was warmed up yesterday (Nov.29). We started the cooldown today.
crvcrios@naic.eduMonday, December 6, 2010 13:57:33Reciever is been warmed to purge and do vac. to the dewar. This is do to the power failure of last week.The reciever will be ready for dec-9-2010 observation.
crvcriosWednesday, December 8, 2010 13:47:04The cryo pump was cleaned and purged. The dewar vac. was o.k. and the system started cooldown.
crvcrios@naic.eduThursday, December 9, 2010 10:06:49System is cold and the vac valve was closed.
crvTuesday, March 1, 2011 11:48:06The cryopump was replaced on friday 2/28/2011. ED 018 was installed system is cold and vac.pump was removed today.
criosMonday, May 23, 2011 16:16:19The system was sounding contaminated the dewar was partially warmed up and purged. The system is already cold.
criosMonday, June 6, 2011 14:00:56On friday june the 3rd a hybrid was installed on the feed horn . The Rf cables from the the feed were broken on the base and were re-soldered.
Carlos Rioscrios@naic.eduThursday, February 2, 2012 10:15:49 The cooling system had a leak and a flexible line on the transmitter room was replaced. After replacing the line the XB and 327 Mhz were contaminated. The system was turned off on Feb. 1st 2012. Started cooldown today.
Carlos Rios crios@naic.eduThursday, February 2, 2012 10:25:49During mid Dic. 2011 a hybrid was connected to the 327 feed antena. the connection is still in place.
crioscrios@naic.eduMonday, May 7, 2012 10:14:17The cryo pump sounded contaminated. The system was turned off.
Denis Urbaindurbain@naic.eduWednesday, April 3, 2013 13:15:57Compressor lost pressure during the weekend. The receiver was warming up because of the lost cooling capacity. The cold head was stopped to restore the vacuum. After 12 hours the vacuum was good and the cold head started again. Until now the compressor has kept is pressure, no clear explanation for the loss of pressure during the weekend.
Denis Urbaindurbain@naic.eduThursday, June 6, 2013 08:08:22The compressor failed because the fan stopped working and the over-temperature sensor was diconnected. The phase monitor, the base of the timer and some fuses had to be replaced. The cold head was cleaned and the vacuum restored before cooling the receiver again on the 6/5/2013.
Felix O. Fernandezfelix.fernandez@naic.eduWednesday, February 19, 2014 07:27:35The cold-head was having some ratcheting sounds on 14-FEB-2014. Based on the intermittent sound the contamination seemd to be minimal. It was decided to have the cold-head warm up and replaced. The replacement took place on 15-FEB-2104.
Severe Noise on Pol Adwhitlow@naic.eduWednesday, May 28, 2014 09:01:19Chris & Tapasi reported a much much higher than normal "noise" seen on Pol A of the 327 MHz RX. Investigation with a square law detector confirms this, but it's not noise. it is a rectangular upwards pulse series, with the pulse duration being a bit over 100 msec and the pulse period being a bit over 200 msec most of the time. The pulse period is irregular, however, with occasional gaps between pulses reaching several times the norm. "Covering" the feed makes the interference go away. This has been found to be due to a pair of closely-spaced carriers in the passband, either of which show the waveform when examined by itself. Phil recorded some of the broadband data with the Mocks, but we have not yet taken a closer look due to time pressure from other urgent projects. We do not yet know if this signal arises from the observatory or from some outside source.
Higher than normal Tsysfelix.fernandez@naic.eduMonday, August 4, 2014 13:45:26The power supply for the Jenning switches that control the ANT/LOAD state blew a fuse last Friday AUG-1-2014. The 327 receiver and Pol-A from the 430G receiver appeared to be warm during this past weekend’s Tsys. The fuse was replaced today and all systems are back to normal.
Polarization changegrivera@naic.eduFriday, August 8, 2014 15:00:51The hybrid was removed from the antenna, per user request.
Polarization hybrid in placefelix.fernandez@naic.eduFriday, September 26, 2014 11:27:13The Hybrid (placed between the feed and the filters) was inserted again into the system today. The Tsys without the hybrid as 86.2/88.7 and now with the hybrid the Tsys is 94.9/104.8
Cold-head replacementfelix.fernandez@naic.eduFriday, May 29, 2015 13:57:22The 20K stage started showing temp oscillations. These oscillations are a typical sign of a worn out seal on the second stage displacer. The cold-head was replaced and a thermal cycle was completed.
Helium leakfelix.fernandez@naic.eduThursday, October 15, 2015 21:33:37Went up to the Dome tonight and found the supply preassure to be ~250psi. Since we recharged the system two days ago it seems the compressor is loosing about 20psi per day. We need to correct this leak.
Helium leak repairfelix.fernandez@naic.eduFriday, October 16, 2015 11:08:45Today we found a leak at the fitting connected to the supply port of the compressor. There was a damaged o-ring that was replaced and the leak was fixed. During the repair 327 lost its vacuum. Hence, it was decided to start a thermal cycle and start cooling 327 ASAP.
Helium leakfelix.fernandez@naic.eduMonday, October 19, 2015 09:50:31The helium leak was still present over the weekend. Today we found a larger leak at the cold-head's motor. It seems there is a faulty o-ring that needs to be replaced. Due to the current observations we plan to complete this procedure on Thursday.
Compressor failurefelix.fernandez@naic.eduWednesday, December 2, 2015 14:53:16Early on Wed Nov-25 the 327/XB compressor failed. After replacing a bad fan motor capacitor, the compressor was placed back online on Monday Nov-30.
PolA - Stage #2 -Idfelix.fernandez@naic.eduThursday, March 30, 2017 11:35:52After the HF campaign (MAR/13-MAR/20) PolA's stage #2 Id became erratic. The Id was "beating" at 1.2Hz, which is the cold-head's frequency of rotation. We tapped on the dewar and the gain instability (due to improper biasing) correlated with the strikes. After searching for possible causes related to anything (i.e. chassis, cables, etc.) on the outside we warmed up the receiver on MAR/28. On MAR/29 we opened the dewar and there was no clear evidence of anything wrong inside the dewar. After reconnecting the bias cable the problem got resolved. We deiced to close the dewar and cool down the system. The system is now (MAR/30) back into operational status. The most likely cause for this bias instability seem to be a bad connection at the LNA. This has been observed in other LNAs which share the same type of connectors (round 7 pin Microtech) which use Teflon within their physical arrangement. These tend to become "loose" at cryogenics temperatures after a prolonged use. A possible final solution would be to replace the bias connectors at the LNAs with D-Sub Micro-D Connectors.


Back to Top